Rogue Fitness equipment produces barbells that are sold in a highly competitive market. Currently, barbells sell for $20 per barbell. Rogue Fitness' total and marginal cost curves are:

            TC = 9,000 + 0.1Q2

            MC = 0.2Q

Calculate Rogue Fitness' profit maximizing quantity. Is the firm earning a profit at the profit maximizing quantity?
Answer #1

Answer : For highly competitive market the profit maximizing condition is P = MC. So,

20 = 0.2Q

=> Q = 20 / 0.2

=> Q = 100

TR (Total revenue) = P * Q = 20 * 100 = $2,000.

TC = 9,000 + 0.1 (100)^2

=> TC = $10,000

Profit = TR - TC = 2,000 - 10,000 = - $8,000

Therefore, here at profit-maximizing quantity level the firm is not earning profit.